log12 (1/144)
I know that because it is 1/144 the answer will be negative but I'm not really sure about anything else. If you could explain to me how logs work that'd be great! Thank you.

Answer by lwsshak3(6460)

log12 (1/144)
I know that because it is 1/144 the answer will be negative but I'm not really sure about anything else. If you could explain to me how logs work that'd be great! Thank you.
..
One definition of a logarithm I find most useful is this:
The base raised to the logarithm of a number=number. This is called the exponential form of a logarithm.
In your case, the base=12, the number is 1/144 and you are trying to find the logarithm of 1/144.
In other words, what do you raise 12, to equal 1/144.
12^x=1/144
1/144=1/12^2=12^-2
12^x=12^-2
since the bases are equal, the exponents must be equal to each other, that is,
x=-2, the logarithm of 1/144, base 12.
To check: raise 12 to -2=12^-2=1/12^2=1/144